Suppose a monopolist faces the following demand curve: P=250-Q

Marginal cost of production is constant and equal to $10, there are no fixed costs

What is the monopolist's profit-maximizing level of output?
Answer #1

Demand curve = 250-Q

Marginal revenue = twice the slope of the demand curve so MR = 250-2Q

MC = 10

Set MC=MR

250-2Q=10

250-10 = 2Q

Q = 240/2 = 120